OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. D(1, 2) and E(–3, 6)

OpenStudy (anonymous):

please someone help me :(

OpenStudy (aum):

Midpoint is: \(\Large (\frac{x_1+x_2}{2}, \frac{y_1+y_2}{2})\)

OpenStudy (anonymous):

meaning???

OpenStudy (aum):

Midpoint is: \(\Large (\frac{1 + (-3)}{2}, \frac{2+6}{2}) \) Simplify.

OpenStudy (aum):

Add the x-coordinates of the end points and divide it by 2. Add the y-coordinates of the end points and divide it by 2. That will be the coordinates of the midpoint.

OpenStudy (anonymous):

And that would come out to be???

OpenStudy (aum):

Can't give out answers.

OpenStudy (anonymous):

why?

OpenStudy (aum):

I am here to help people learn math. Not to do their homework or test.

OpenStudy (anonymous):

its not for home work or a test :|

OpenStudy (aum):

You just need to add two numbers and divide by 2 and you should be able to do that.